0 0
Add a comment Improve this question Transcribed image text
Answer #1

Use the activity formula,

T1/2 = 4.468 x 10^9 yr x 365 x 24 x 3600 = 1.409 x 10^17 sec

l = ln(2)/T1/2 = ln(2)/1.409 x 10^17 = 4.91932697 x 10^-18 s-1

DN/Dt = lN,

or, 730 = 4.91932697 x 10^-18 x N

N = 1.48 x 10^20 nuclei (answer)